Hours of sitting per day significantly predicts the amount of weight gain (in 6 months). The regression equation is WeightGain = 7 + 0.49 * HoursSit. Data are from 120 full-time employees in a public urban university. With a standard error for the slope (SE_b1) of 0.02, what is the 95% confidence interval of the slope.

95%CI:(6.9608, 7.0392), which means the slope is statistically significant.

95%CI:(0.4508, 0.5292), which means the slope is statistically significant.

95%CI:(6.9608, 7.0392), which means the slope isn't statistically significant.

95%CI:(0.4508, 0.5292), which means the slope isn't statistically significant.

need help answering question
Answer #1

n=120, s.e. 0.02, t=1.98(approx),slope=b1=0.49

upper interval=0.49+0.02*1.98=0.5296

lower interval=0.49-0.02*1.98=0.4504

the answer i evaluated is not exactly equal to the answer of the mcq ques but it's approximately equal and the error is negligable

conclusion:

95%CI:(0.4508, 0.5292), which means the slope isn't statistically significant.